Solve the following equations:

(a) 5x + 2 = 7 (b)

6x + 5 = 3x + 11

(c) 2x - (1-3x) = 8x - 2(3x - 4)
Answer #1

(a) Curen can : 5x + 2 = 7. To elinunate 2 from 2.H.S , we add (-2) to both sides of the ear. 50+2 +(-2) = 7+ (-2). or 50+ 2-. $ (34) = (6) on as 2 Thus the solution is: x=2. x=2. An (e) 2x - (1-3) = 8x-2 (3-4) 2x-1 +3x = 8x - 6x + 8 50-1 a 20 + 8 0